How many atoms are in a spoonful of water?



1
Expert's answer
2021-03-11T08:48:34-0500

We will use 5 mL as the volume of spoonful.

Density of water = 1 g/mL

m = 5 g

M(water) = 18 g/mol

n(water)"= \frac{m}{M} = \frac{5}{18}= 0.278 \;mol"

A mole of water contains 6.02×1023 molecules.

"N_{molecules} = 0.278 \times 6.02 \times 10^{23} = 1.67 \times 10^{23}"

But each water molecule is comprised of 3 atoms: two hydrogen and one oxygen.

So, the total number of atoms is "3 \times 1.67 \times 10^{23} = 5.01 \times 10^{23}"